OpenStudy (anonymous):

6x-2y=33 4x+3y=9 Solve the simultaneous equation

myininaya (myininaya):

I would try elimination. First I multiply first equation by 3 and second equation by 2. \[3(6x-2y=33)=>18x-6y=99\] \[2(4x+3y=9)=>8x+6y=18\] -------------- now add equations -> 26x+0=117 -> 26x=117 -> x=117/26 Now plug this into one of equations given to you to find y

OpenStudy (anonymous):

multiply first equation with 3/2 it became 9x-3y=99/2 second is 4x+3y=9 add them 13x=117/2 x=117/26 if you put x to the each equations you will easily find y